### v2

@@ B.4 Extended representational analysis
By prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022),
@@
Appendix C Detailed methods
C.1 Models
@@
We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn Pedregosa et al. (2011). Note that this is equivalent to using a feature basis with the same resulting similarities. However, hand-specifying these similarities enabled us to easily explore the full range of possible representations.
@@
We considered networks with four convolutional layers (kernel size is five, two layers have 32 filters, two have 64 filters) and two densely connected layers (with 512 and 1024 units). Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation. All weights are initialized with He initialization (He et al., 2015).
@@ C.2 Additivity analysis
Appendix D Compositional task space
D.1 Symbolic addition
D.1.1 Proof of Proposition 5.2
See 5.2
@@ D.1 Symbolic addition
D.1.2 Behavior of deep ReLU networks
@@
D.1.3 Behavior of vision models
@@
D.2 Context dependence
D.2.1 General task definition
@@
D.2.2 Novel stimulus compositions are conjunction-wise additive
@@
D.2.3 Novel rule compositions are not conjunction-wise additive
@@ D.2.4 Coefficient groups
D.2.5 Feature-learning ReLU networks
We find that feature-learning ReLU networks generalize consistently on CD-1 anREGULARIZATIONd CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 10).
@@ D.2.6 Vision networks
D.3 Invariance and partial exposure
@@
D.4 Other mathematical operations
@@ D.5 Logical operations
D.6 Transitive ordering
Transitive ordering is a popular task in cognitive science (often called transitive inference, (McGonigle & Chalmers, 1977)). Here the subject is presented with two items $z_{1}$, $z_{2}$ drawn from an unobserved hierarchy $>$. It should then categorize whether $z_{1}>z_{2}$ or $z_{2}>z_{1}$. Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ (Lippl et al., 2024). It is therefore additive, in contrast to transitive equivalence. This is also the case if we assume that there are multiple such hierarchies (e.g. $a_{1}>\dotsc,a_{5}$ and $b_{1}>\dotsc,b_{5}$). In this case, the model would generalize to comparisons between these different hierarchies as well.
Appendix E A feature-learning mechanism for non-additive compositional generalization
@@
We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence. To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input. By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime. Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs. 7 and 10). On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig. 13a).

### v3

@@ B.5 Salience in very deep ReLU networks
By prior work (Cho & Saul, 2009; Tsuchida et al., 2018; 2019; Han et al., 2022),
@@
Appendix C DSprites representations
To investigate whether our theory can describe compositional generalization in practically used disentangled models, we considered six different model architectures considered in Locatello et al. (2019) and trained on the DSprites dataset, an important benchmark for disentangled representation learning. This paper considers fifty random seeds for each model and further considers six different possible hyperparameters per architectures, resulting in a total of 1,800 models. In our analysis, we only analyze differences between architectures, pooling across all hyperparameter choices.
C.1 Task setup
DSprites consists of small black-and-white shapes and has five different underlying components: shape (three categories: hearts, ovals, and squares), size, x-position, y-position, and rotation. We consistently consider the largest possible size and hold the rotation fixed at zero. While x- and y-position can each take on 32 possible values, we only consider four different categories for each. Importantly, this is still a highly non-compositionally structured representation: not only do the disentangled representation learning methods often fail to discover the underlying factors of variation; the model also likely represents x- and y-positions that are closer to each other, as more similar, violating our compositionally structured assumption. These representations therefore present a particularly challenging test of our framework.
For symbolic addition and transitive equivalence, we consider x- and y-position as the two components. For each task instance, we randomly determine which position takes on which component’s role. In total, we consider 50 randomly sampled task instance for each of the 1,800 models. For context dependence, we subsample two shapes and assume that these shapes provide the context cue. We then consider x- and y-position as feature 1 and feature 2. While the context dependence considered in the main text had six possible feature values for each feature, we now only consider four. As a result, we only consider two different datasets: CD-2, which leaves out all trials where feature 1 indicates category 2 and feature 2 indicates category 1; and CD-1 which only leaves out one conjunction of features. Put differently, on a given trial, the task could for example look as follows: if we see a heart shape, the model should output whether this object is in a certain x-position; if we see a square shape, the model should output whether this object is in a certain y-position. While the model has seen all x- and y-positions individually, it has not seen each combination of x- and y-positions.
Finally, we consider either a direct linear readout from the disentangled representation or an initial transformation by a one-hidden-layer ReLU neural network with random weights.
Figure 8: Conjunction-wise additivity in empirical disentangled models. a, Additivity across the different model architectures and four different task settings. All of the model behaviors are generally highly additive. b, Distribution of accuracies on transitive equivalence when averaged across different component roles. c, Variance ratio, capturing the ratio between the deviation between the representational similarity matrix and its compositionally structured approximation and the general variance of the representational similarity matrix. The numbers indicate that all disentangled models discover some compositional structure in their representation, but still deviate substantially from a compositionally structured representation. d, Generalization error on symbolic addition plotted against the estimated $S(1;2)$. e, Average accuracy on two versions of context dependence with a total of four features: CD-2 which now leaves out an entire orthant, and CD-1, which only leaves out a single data point. Again, models generally perform below chance on CD-2 but above chance on CD-1, though performance is generally really low.
C.2 The models are well described by a conjunction-wise additive computation
We first investigated whether the model predictions on the test set were well characterized by a conjunction-wise additive computation (see Section D.3), averaging this model behavior across the fifty random task instances. We found that they were generally well captured, though they were certainly not perfectly conjunction-wise additive (Fig. 8a). Further, none of these models were able to systematically generalize on transitive equivalence (Fig. 8b). This indicates that conjunction-wise additivity may characterize the generalization class of these highly non-compositionally structured representations as well — at least when averaged across task instances.
C.3 The representations are partially compositionally structured
We next investigated whether these representations exhibit compositional structure according to our definition. To determine this, we computed their representational similarity matrix and computed the average similarity for each set of overlaps. This instantiates the compositionally structured representational similarity matrix that most closely described the empirical representational similarity. We then determined the average squared deviation from this matrix, $\sigma_{cs}^{2}$, comparing it to the overall variance of this matrix, $\sigma_{var}^{2}$. Overall, the variance ratio, $\sigma_{cs}^{2}/\sigma_{var}^{2}$, characterizes the degree to which the given representation is compositionally structured. In a fully non-compositionally structured representation, the variance ratio will be roughly one, whereas in a fully compositionally structured representation, it will be zero. For the disentangled representation learning models, we found that this variance ratio took on value between 0.5 and 0.9 (Fig. 8c). Thus, the models’ representation partially captured compositional structure, but also contained a lot of non-compositional structure.
C.4 The non-compositional representational structure substantially impacts model behavior
Finally, we tested whether our representational geometry analysis in Section 5 extended to the DSprites representations. We first estimated their salience $S(1;2)$ by computing the average representational similarity for each overlap and then computing the salience using Eq. 13. Proposition 5.1 would predict that the generalization error should decrease with increasing $S(1;2)$. However, we do not observe such a trend (Fig. 8d). This indicates that our representational geometry analysis does not apply to the DSprites representations, due to the way in which they violate the compositional structure assumption.
Finally, we determined generalization accuracy on CD-2 and CD-1. We found that most models performed quite poorly, perhaps owing to the fact that shape is barely represented in these models. Nevertheless, we found that most models performed systematically below chance for CD-2 but systematically above chance for CD-1. Our analysis of context dependence in compositionally structured representations suggests that this may be because these models exploit a context-driven shortcut on CD-2 but not CD-1.
Overall, our analysis therefore paints a nuanced picture of the applicability of our theory to practical disentangled representations. Our findings suggest that these models may still be well approximated by conjunction-wise additive computations and that this generalization class may therefore shed light on fundamental limits to the generalization of linear readout models. At the same time, to understand how these models generalize on conjunction-wise additive tasks, taking into account the specific ways in which they violate the compositional structure assumption is likely important.
Appendix D Detailed methods
D.1 Models
@@
We fit the kernel models by hand-specifying the kernel and fitting either a support vector regression or classification using scikit-learn (Pedregosa et al., 2011). Note that this is equivalent to using a feature basis with the same resulting similarities. However, hand-specifying these similarities enabled us to easily explore the full range of possible representations.
@@
Compositional MNIST/CIFAR-10.
We create a compositional version of MNIST and CIFAR-10 by concatenating multiple images along different channels. For example, the input to the MNIST version of the symbolic addition task had two channels, each containing one digit whereas the CIFAR-10 version of the symbolic addition task had six channels, three for each digit. We further varied the distance between the concatenated images either presenting them all on top of each other or presenting them offset by a certain number of pixels. Each image category corresponded to a certain component, where its role as ranodmly sampled for each task instance. The output was still given by a single scalar: the total magnitude for symbolic addition and the two categories for context dependence and transitive equivalence.
@@
We considered networks with four convolutional layers (kernel size is five, two layers have 32 filters, two have 64 filters) and two densely connected layers (with 512 and 1024 units). Each layer is followed by a ReLU nonlinearity, and the convolutional stage is followed by a max pooling operation. All weights are initialized with He initialization (He et al., 2015). We trained these networks with SGD using a learning rate of $10^{-4}$ and momentum of $0.9$.
@@ D.3 Additivity analysis
Appendix E Compositional task space
E.1 Symbolic addition
E.1.1 Proof of Proposition 5.1
See 5.1
@@ E.1 Symbolic addition
E.1.2 Behavior of deep ReLU networks
@@
E.1.3 Behavior of vision models
@@
Overall, our analysis indicates that our theory provides several valuable qualitative insights into the behavior of deep neural network models. Importantly, however, we are not able to provide quantitative bounds at the moment. To demonstrate this, we computed the salience $S(1;2)$ of the neural tangent kernel of the convolutional neural networks before and after being trained on the MNIST version of symbolic addition with $\mathcal{W}=\{0\}$ (Fig. 13a). Importantly, $S(1;2)$ changed substantially during training, indicating that these networks are not trained in the kernel regime. As such, our theory does not apply exactly. To see whether it provided appropriate quantitative bounds, we then estimated the slope of each model instance on the compositional generalization dataset and plotted it against $S(1;2)$ of the neural tangent kernel at the beginning of training. We found that higher $S(1;2)$ generally resulted in a larger slope, as predicted by our theory. However, the relationship between $S(1;2)$ and slope did not adhere exactly to the quantitative predictions made by our theory. This indicates that while our theory can shed light on certain qualitative behaviors, it still leaves certain generalization behaviors in deep neural networks unclear.
Figure 13: a, $S(1;2)$ in the convolutional neural network’s neural tangent kernel, before and after being trained on symbolic addition with $\mathcal{W}=\{0\}$. b, The estimated slope for each model instance plotted against $S(1;2)$. While increased $S(1;2)$ generally yields a larger slope (as predicted by our theory), the relationship between $S(1;2)$ and the slope does not follow the exact quantitative relationship predicted by our theory.
Finally, our analysis in Section 5.1 suggests that deeper models should yield more conjunctive representations, and therefore exhibit worse generalization on symbolic addition. To test this prediction, we varied the number of fully connected layers in the convolutional neural network and trained these networks on the symbolic addition task with the training set $\mathcal{W}=\{0\}$. We found that deeper networks indeed generalized worse (Fig. 14).
Figure 14: Mean squared error on the compositional generalization set for different depths of the fully connected network at the end of the convolutional neural network architecture.
E.2 Context dependence
E.2.1 General task definition
@@
E.2.2 Novel stimulus compositions are conjunction-wise additive
@@
E.2.3 Novel rule compositions are not conjunction-wise additive
@@ E.2.4 Coefficient groups
E.2.5 Feature-learning ReLU networks
We find that feature-learning ReLU networks generalize consistently on CD-1 and CD-2 but not CD-3. They are also perfectly conjunction-wise additive and fail due to a context shortcut (Fig. 15).
@@ E.2.6 Vision networks
E.3 Invariance and partial exposure
@@
E.4 Other mathematical operations
@@ E.5 Logical operations
E.6 Transitive ordering
Transitive ordering is a popular task in cognitive science (often called transitive inference, McGonigle & Chalmers (1977)). Here the subject is presented with two items $z_{1}$, $z_{2}$ drawn from an unobserved hierarchy $>$. It should then categorize whether $z_{1}>z_{2}$ or $z_{2}>z_{1}$. Crucially, this task can be solved by assigning a rank $r(z_{c})$ to each item and computing the response as $f(z)=r(z_{1})-r(z_{2})$ Lippl et al. (2024). It is therefore additive, in contrast to transitive equivalence. This is also the case if we assume that there are multiple such hierarchies (e.g. $a_{1}>\dotsc,a_{5}$ and $b_{1}>\dotsc,b_{5}$). In this case, the model would generalize to comparisons between these different hierarchies as well.
Appendix F A feature-learning mechanism for non-additive compositional generalization
@@
We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence. To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input. By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime. Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs. 10 and 15). On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig. 18a).
@@
Figure 19: This figure composes the relevant theoretical insights and empirical experiments to illustrate how they speak to each other.